OpenStudy (anonymous):

Evaluate the indefinite integral of: 1 ---------- cos(x) - 1 dx I would like a hint please, not the answer.

OpenStudy (turingtest):

\[\int{dx\over\cos x-1}\]? that's tricky I think...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what if u miltiply it by (cosx+1)?

OpenStudy (turingtest):

that should do it!

OpenStudy (turingtest):

just be be clear @QRAwarrior shinigami meant multiply it by\[{1+\cos x\over1+\cos x} \]and use some trig identities that is a sufficiently good tip

OpenStudy (anonymous):

rationalization? am i right?

OpenStudy (anonymous):

But then how would U-subsitution work from there on?

OpenStudy (anonymous):

do you know half angle method?

OpenStudy (anonymous):

How would you further go on? I get this: 1 + cosx -------- (sinx)^2 dx

OpenStudy (anonymous):

Yes that would be (sinx)^2 = [1/2][1-cos(2x)]

OpenStudy (anonymous):

But U-substitution would still not work...

OpenStudy (anonymous):

its 1+cos/ -sin^2

OpenStudy (anonymous):

Yes I realized.

OpenStudy (anonymous):

But I do not see what I can do with 1 + cos/(1/2)(1 + cos(2x))

OpenStudy (anonymous):

just... divide it dont simplify...

OpenStudy (anonymous):

Got it.